OpenStudy (anonymous):

well fan and medal Which of the following is an equivalent representation of 2^-3 ?

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (anonymous):

i need help plzz!!!

OpenStudy (studygurl14):

\(\large 2^{-3}=\frac{1}{2^3}=\frac{1}{(2)(2)(2)}=\frac{1}{4(2)}=\frac{1}{8}\)

OpenStudy (anonymous):



OpenStudy (anonymous):

i got it right thank you

OpenStudy (studygurl14):

You're welcome. :)